# On Continuous Full-Order Integral-Terminal Sliding Mode Control with Unknown A Priori Bound on Uncertainty

### v3

@@ 4 Chattering free adaptive SMC with unknown apriori bound on uncertainty
$u_{\text{ado}}=$ $-\kappa s-\hat{d},$ (26)
and the disturbance observer dynamics
$\hat{d}=$ $\lambda\left(x_{n}-\zeta\right),$ (27)
$\dot{\zeta}=$ $f_{n}(\textbf{x},t)+b(\textbf{x},t)u+\hat{d}-\frac{\hat{k}}{\lambda}\text{sgn}(\tilde{d})-\frac{s}{\lambda},$ (28)
with an ultimate bound on $||\mathcal{S}||^{2}$ as
$\mathcal{B}=\sqrt{\frac{2\bar{\delta}}{\gamma-\theta}},$
and the adaptation law
$\dot{\hat{k}}$ $=-\tau\hat{k}+\mu|s|,\;\;\hat{k}(0)>0,$ (29)
where $\mathcal{S}=\left[s,\;\tilde{d},\;\tilde{k}\right]^{T}$, $z(0)=x_{n}(0)$, $s(0)=s_{0}$, $\hat{k}(0)\in\mathbb{R}^{>0}$, $\tau=\left(\mu+1+\tau_{0}\right)$, $\lambda,\mu,\tau_{0},\kappa\in\mathbb{R}^{+}$, $\bar{\delta}=\frac{1}{2}\tau k^{2}$. The error between actual ($d$) and estimated disturbance ($\hat{d}$) is given by
$\tilde{d}=d-\hat{d},$
and, the error etween the actual $k$ and an estimation of $k$ as $\hat{k}$ is given by
$\tilde{k}$ $=k-\hat{k}.$
@@
The governing disturbance observer dynamics in (27) and (28), a dynamical equation in $\zeta$, is continuous since the discontinuity lies in $\dot{\zeta}$.
@@
Fig. 2 presents a schematic diagram illustrating the proposed control algorithm. The disturbance observer in (28) reconstructs the disturbance considering the plant states, the control signal, the adaptive gains and the sliding variable and feeding them to the controller. The controller in (26) uses a state feedback mechanism also making use of the sliding variable, estimated disturbance and the adaptive gain data manufactured using adaptive laws. This architecture ensures good performance in the presence of matched disturbance of bounded nature but the bounds being unknown.
@@
The Lyapunov stability analysis of the closed-loop system uniquely incorporates the dynamics of disturbance error, which sets it apart from other designs of disturbance observers found in existing literature.
@@
In eq (28), a term containing $\text{sgn}(\tilde{d})$ has been used. Since $d$ is not available for measurement, we define a function,
$\omega(t)=$ $\int_{t_{0}}^{t}\tilde{d}(\sigma)\mathrm{d}\sigma,$
$=$ $x_{n}-\int_{t_{0}}^{t}\left(f(\mathbb{x},\sigma)+b(\mathbb{x},\sigma)u(\sigma)+\hat{d}(\mathbb{x},\sigma)\right)\mathrm{d}\sigma,$
and define $\text{sgn}(\tilde{d})$ as $\text{sgn}(\tilde{d}(t))=\text{sgn}\left(w(t)-w(t-\tau_{d})\right)$ where $\tau_{d}$ is the time delay.
Since $\tilde{d}=\underset{t\to 0}{\text{lim}}\frac{\left(w(t)-w(t-\tau_{d})\right)}{\tau_{d}}$, one can choose $\tau_{d}$ as fundamental sampling time of the system. It is to be noted we need to know the sign of $\tilde{d}$ instead of $\tilde{d}$ whose calculation is comparatively easy.
@@
The dynamical equation (28) constitutes the disturbance observer ([38, 39]) for the system (5) with the characterisation of disturbance $d$ as given in (23).
5 Illustrative Example II
@@
where the disturbance $d=\text{sin}(2\pi t)$. The sliding surface is chosen the same as (22). The control input $u$ is designed as,
@@
(a) States with adaptive ADO based FOITSMC
(b) Sliding variable of adaptive ADO based FOITSMC
(c) Input for FOITSMC with adaptive ADO
(d) Disturbance estimation error for adaptive ADO based FOITSMC
(e) Estimation of $k$ for FOITSMC with adaptive ADO
Figure 3: 3a, 3b and 3d show the ultimate boundedness of state trajectories, sliding variable and disturbance estimation error, respectively, on the application of adaptive ADO-based FOITSMC. Moreover, the minimum estimated adaptive gain required for GUUB is shown in 3e.
@@
The strategies employed in adaptive STC [26, 31, 34, 35] involve signum functions in adaptation laws, which may potentially lead to unnecessary overestimated (underestimated) gains. This can occur when the gain value increases even as the magnitude of the sliding variable decreases (or increases). In contrast, the adaptive law in (29) is devoid of any such discontinuous functions and strives to attain the least possible value of $\hat{k}$ to obtain the desired result.
These simulation results are obtained using the ODE 4 (Runge-Kutta) solver of MATLAB R2021a with a fixed step size of 0.001. In fig. (3a) and fig. (3b), the states of (30) and sliding variable, respectively, are circumscribed within a specified bound in finite time on the implementation of input in eq. (31). Fig. (3c) displays a control input free from any chattering phenomenon. The estimated $k$ is non-monotonic (as shown in fig. 3e) and it decreases (or increases) depending on whether the magnitude of the sliding variable decreases (or increases).
6 Performance Comparison
@@
such that $|\dot{a}_{1}(x,t)|<\delta<\infty$, $\delta$ unknown. For any given arbitrary initial conditions $x(0)$, $\sigma(0)$, there exists a finite $t_{f}>0$ such that a real $2-$sliding mode has been established $\forall\;t>t_{f}$ via control law
@@
$\dot{\alpha}$ =$=\begin{cases}\omega_{1}\sqrt{\frac{\gamma_{1}}{2}}\text{sgn}\left(|\sigma|-\mu\right),\;\;&\alpha>\alpha_{m}\\ \eta,\;&\alpha\leqslant\alpha_{m}\end{cases}$ (34)
@@
(a) The adaptive gain $(\hat{k})$ for disturbance
$d=\sin(2\pi t)$
(b) The adaptive gain $(\hat{k})$ for disturbance
$d=\sin(2\pi t)+\texttt{ramp}(t)$
(c) The sliding variable $(s)$ in presence of
disturbance $d=\sin(2\pi t)$
(d) The sliding variable $(s)$ in presence of
disturbance $d=\sin(2\pi t)+\texttt{ramp}(t)$
[…]
7 Case study
The attitude control and stabilization of a rigid spacecraft in the presence of external disturbances have been a topic of interest for control enthusiasts for quite a long time ([41, 42, 43, 44, 45, 46]) because of its highly nonlinear dynamical nature which makes the control algorithm design much more challenging and persevering. This section portrays an application of the aforementioned control algorithm for the attitude stabilization of a rigid spacecraft. The unit-quaternion-based attitude dynamics of a rigid spacecraft are given by
$\begin{cases}\dot{q}_{0}&=-\frac{1}{2}q_{v}^{T}\Omega\\ \dot{{q}_{v}}&=0.5\left(q_{0}I_{3}+q_{v}^{\times}\right)\Omega\\ J\dot{\Omega}&=-\Omega^{\times}J\Omega+u+d.\end{cases}$
The unit-quaternion is defined by $q=[q_{0}\;q_{1}\;q_{2}\;q_{3}]^{T}=[q_{0}\;q_{v}]^{T}$ satisfying a constraint $q_{0}^{2}+q_{v}^{T}q_{v}=1$, where $q_{v}\in\mathbb{R}^{3}$ is the vector part and $q_{0}\in\mathbb{R}$ is the scalar component, $\omega\in\mathbb{R}^{3}$ is the angular velocity of the spacecraft, $u\in\mathbb{R}^{3}$ is the control torque, $J\in\mathbb{R}^{3\times 3}$ is the symmetric positive definite inertia matrix, $I_{3}$ is the $3\times 3$ identity matrix, and $\times$ is an operator on any vector $a=[a_{1}\;a_{2}\;a_{3}]^{T}$ such that,
$a^{\times}$ $=\begin{pmatrix}0&-a_{3}&a_{2}\\ a_{3}&0&-a_{1}\\ -a_{2}&a_{1}&0\end{pmatrix},$
@@ 7 Case study
The exogenous matched disturbance $d$ is continuously differentiable such that each components $d_{i}$ of $d$ satisfy,
$\dot{d}_{i}\leqslant k_{i},\;\;\text{for}\;i=1,2,3,$
where $k_{i}$s are finite.
@@
$\implies\dot{V}=$ $s^{T}\left(J^{-1}\tilde{d}-\Theta s\right)+\tilde{d}^{T}\dot{d}-\tilde{d}^{T}\Lambda J^{-1}\tilde{d}-\tilde{d}^{T}\hat{k}\text{sgn}(\tilde{d})$
@@
$k_{i}\hat{k}_{i}-\hat{k}_{i}^{2}$ $=-\left(\frac{k_{i}}{\sqrt{2}}-\frac{\hat{k}_{i}}{\sqrt{2}}\right)^{2}-\frac{\hat{k}_{i}^{2}}{2}+\frac{k_{i}^{2}}{2},$
$\implies$ $\leqslant-\frac{\tilde{k}_{i}^{2}}{2}+\frac{k_{i}^{2}}{2}.$ (38)
@@
where $\Gamma=\text{min}\left(\lambda_{\text{min}}(\bar{\Theta}),\lambda_{\text{min}}(\bar{\Lambda}),\lambda_{\text{min}}(\frac{\tau_{0}}{2})\right)$, $\bar{\Theta}=\Theta-\frac{\lambda_{\text{max}}(\mu)}{2}I_{3}$, $\bar{\Lambda}=\Lambda J^{-1}-\frac{1}{2}I_{3}$ and $\bar{\Delta}=\lambda_{\text{max}}(\tau)||k||^{2}$. On similar lines to that of eq. (LABEL:pf_3) and (LABEL:lya_guub1), eq (39) can be expressed as
@@
(a) Evolution of quaternions
(b) Evolution of angular velocities
(c) Sliding variable of FOITSM for rigid spacecraft dynamics
(d) Disturbance estimation error of adaptive ADO with FOITSMC
(e) Estimation of $k$
(f) Input for FOITSMC with adaptive ADO
Figure 5: 5a, 5b and 5c show the global boundedness of quaternions, $\Omega$, and sliding variable, respectively, in the presence of matched disturbances. 5d shows the global boundedness of disturbance estimation error. The continuous control using adaptive ADO-based FOITSMC and the estimated value of $k$ is shown in 5e and 5f, respectively.
@@
The simulation results shown in fig. 5a and 5b, utilising the ODE 4 (Runge-Kutta) solver of MATLAB R2021a with an assigned fixed step size of 0.001, depict that the state trajectories remain close to the origin, bounded within a specified limit. A continuous control input has been employed using an adaptive ADO-based FOITSMC algorithm, as shown in Fig. 5f, to ensure the global boundedness of error in the estimation of conditionally known disturbance and unknown gain $k$.
8 Conclusion and outlook
This paper presents an algorithm ensuring the Global Uniform Ultimate Boundedness (GUUB) of state trajectories in the presence of disturbances with unknown bounds. Additionally, a sliding manifold is devised that combines the attributes of Integral Terminal Sliding Mode Control (ITSMC) and nonsingular Terminal Sliding Mode Control (TSMC). The Lyapunov-based stability analysis incorporates the dynamics of the Adaptive Disturbance Observer (ADO). Consequently, the monotonically increasing nature of the adaptive gain is eliminated. For future work, exploring the effectiveness of the algorithm for dynamical systems that are not affine in control input, designing an observer-based controller for the same, modifying the algorithm to work in the realm of mismatched perturbations, and determining a more stringent bound on states and errors can be a potential research objective.
Disclosure statement
The authors have no conflict of interest to disclose.
Appendix
8.1 Proof of Theorem 2.3
Differentiating (6) once with respect to time
$\dot{s}=$ $\dot{x}_{n}+C_{n}|x_{n}|^{\alpha_{n}}\text{sgn}(x_{n})+C_{n-1}|x_{n-1}|^{\alpha_{n-1}}\text{sgn}(x_{n-1})$
$+...+C_{1}|x_{1}|^{\alpha_{1}}\text{sgn}(x_{1}).$
Substituting the dynamics of $x_{n}$ from (5)
$\dot{s}$ $=f_{n}(\textbf{x},t)+b(\textbf{x},t)u+d(\textbf{x},t)+C_{n}|x_{n}|^{\alpha_{n}}\text{sgn}(x_{n})$
$+C_{n-1}|x_{n-1}|^{\alpha_{n-1}}\text{sgn}(x_{n-1})+...+C_{1}|x_{1}|^{\alpha_{1}}\text{sgn}(x_{1}).$
Substituting $u$, $u_{eqv}$, and $u_{dis}$, respectively, from (17), (18) and (19),
$\dot{s}$ $=d(\textbf{x},t)+f_{\Delta}(\textbf{x},t)-(\eta+d_{max}+f_{max})\text{sgn}(s),$
$\leqslant-\eta\;\text{sgn}(s),$
which implies that $s$, if it was initially at zero, remains at zero for all time $t\geqslant 0$. Hence, the system evolves by following the dynamical equation (16) from the initial time $t=0$. This completes the proof.
8.2 Proof of Theorem 4.1
Consider a Lyapunov function candidate as
$V=$ $\frac{1}{2}s^{2}+\frac{1}{2}\tilde{d}^{2}+\frac{1}{2}\tilde{k}^{2},$
$\dot{V}=$ $s\left(\tilde{d}-\kappa s\right)+\tilde{d}\left(\dot{d}-\dot{\hat{d}}\right)+\tilde{k}(-\dot{\hat{k}}),$
$=$ $-\kappa s^{2}+s\tilde{d}+\tilde{d}\left(\dot{d}-\lambda\tilde{d}-\hat{k}\text{sgn}(\tilde{d})-s\right)$
$+\tilde{k}(-\dot{\hat{k}}),$
$=$ $-\left(\kappa-\frac{1}{2}\mu\right)s^{2}-\left(\lambda-\frac{1}{2}\right)\tilde{d}^{2}+\frac{1}{2}(\mu+1)\tilde{k}^{2}+\tau\tilde{k}\hat{k},$
$+\tau\left(k\hat{k}-\hat{k}^{2}\right).$ (40)
Using the fact that $k\hat{k}\geqslant 0$ for all time and,
$k\hat{k}-\hat{k}^{2}$ $=-\left(\frac{k}{\sqrt{2}}-\frac{\hat{k}}{\sqrt{2}}\right)^{2}-\frac{\hat{k}^{2}}{2}+\frac{k^{2}}{2},$
$\implies$ $\leqslant-\frac{1}{2}\tilde{k}^{2}+\frac{k^{2}}{2}.$ (41)
Using (41), (40) becomes
$\dot{V}\leqslant$ $-\left(\kappa-\frac{1}{2}\mu\right)s^{2}-\left(\lambda-\frac{1}{2}\right)\tilde{d}^{2}$
$-\frac{1}{2}\left(\tau-\mu-1\right)\tilde{k}^{2}+\frac{1}{2}\tau k^{2},$
$\implies\dot{V}\leqslant$ $-\gamma V+\bar{\delta},$
For $0<\theta<\gamma$, the above equation can be rewritten as,
$\dot{V}$ $=-\theta V-\left(\gamma-\theta\right)V+\bar{\delta}.$
As the term containing sliding variable $\frac{1}{2}s^{2}\leqslant V$, the state trajectories are also confined within the closed ball of radius $\sqrt{\frac{2\bar{\delta}}{\gamma-\theta}}$. This ends the proof.

### v4

@@ III Chattering free adaptive SMC with unknown apriori bound on uncertainty
$u_{\text{ado}}=$ $-\kappa s-\hat{d}$. (9)
The disturbance observer dynamics are defined as
$\begin{cases}\hat{d}=\lambda(x_{n}-\zeta),\\ \dot{\zeta}=f_{n}(\mathbf{x},t)+b(\mathbf{x},t)u+\hat{d}-\frac{\hat{k}}{\lambda}\operatorname{sgn}(\tilde{d})-\frac{s}{\lambda},\end{cases}$ (10)
with the adaptation law
$\dot{\hat{k}}$ $=-\tau\hat{k}+\mu|s|,\quad\hat{k}(0)>0,$ (11)
where $\tilde{d}=d-\hat{d}$, $\tilde{k}=k-\hat{k}$, $z(0)=x_{n}(0)$, $s(0)=s_{0}$, $\tau=\mu+1+\tau_{0}$, and $\lambda,\mu,\tau_{0},\kappa\in\mathbb{R}^{+}$. For $\mathcal{S}=[s,\tilde{d},\tilde{k}]^{T}$, the ultimate bound on $\|\mathcal{S}\|^{2}$ is:
$\mathcal{B}=\sqrt{\frac{2\bar{\delta}}{\gamma-\theta}},$ (12)
where $\bar{\delta}=\frac{1}{2}\tau k^{2}$ and $\lVert\cdot\rVert$ denotes the usual Euclidean norm.
Proof.
Consider a Lyapunov function candidate as
$V=$ $\frac{1}{2}s^{2}+\frac{1}{2}\tilde{d}^{2}+\frac{1}{2}\tilde{k}^{2},$
[…]
$+\tau\left(k\hat{k}-\hat{k}^{2}\right).$ (13)
Using the fact that $k\hat{k}\geqslant 0$ for all time and,
$k\hat{k}-\hat{k}^{2}$ $=-\left(\frac{k}{\sqrt{2}}-\frac{\hat{k}}{\sqrt{2}}\right)^{2}-\frac{\hat{k}^{2}}{2}+\frac{k^{2}}{2}\leqslant-\frac{1}{2}\tilde{k}^{2}+\frac{k^{2}}{2}.$ (14)
Using (14), (13) becomes
$\dot{V}\leqslant$ $-\left(\kappa-\frac{1}{2}\mu\right)s^{2}-\left(\lambda-\frac{1}{2}\right)\tilde{d}^{2}-\frac{1}{2}\left(\tau-\mu-1\right)\tilde{k}^{2}+\frac{1}{2}\tau k^{2},$
$\leqslant$ $-\gamma V+\bar{\delta},$
For $0<\theta<\gamma$, the above equation can be rewritten as,
$\dot{V}$ $=-\theta V-\left(\gamma-\theta\right)V+\bar{\delta}.$
As the term containing sliding variable $\frac{1}{2}s^{2}\leqslant V$, the state trajectories are also confined within the closed ball of radius $\sqrt{\frac{2\bar{\delta}}{\gamma-\theta}}$. This ends the proof. ∎∎
@@
The governing disturbance observer dynamics in (10), a dynamical equation in $\zeta$, is continuous since the discontinuity lies in $\dot{\zeta}$.
@@
The Lyapunov stability analysis of the closed-loop system distinctively incorporates disturbance error dynamics, which sets it apart from other existing disturbance observer designs in the literature. The schematic in fig. 1 illustrates the control architecture.
@@
[…]
$\omega(t)=\int_{t_{0}}^{t}\tilde{d}(\sigma)\,d\sigma=$ $x_{n}-\int_{t_{0}}^{t}\left(f(\mathbf{x},\sigma)+b(\mathbf{x},\sigma)u(\sigma)\right$.
$\left.\quad+\hat{d}(\mathbf{x},\sigma)\right)\,d\sigma,$ (16)
and define $\text{sgn}(\tilde{d}(t))=\text{sgn}(\omega(t)-\omega(t-\tau_{d}))$, where $\tau_{d}$ is the system’s sampling time. Since $\tilde{d}=\lim_{\tau_{d}\to 0}\frac{\omega(t)-\omega(t-\tau_{d})}{\tau_{d}}$, only the sign of $\tilde{d}$ is needed, which is simpler to compute.
@@
The dynamical equation (10) constitute the disturbance observer for system (5) with the characterisation of disturbance $d$ as given in assumptions 1 and 2.
IV Illustrative Example I
@@
where the disturbance $d=\text{sin}(2\pi t)$. The sliding surface is chosen as follows
$s$ $=x_{3}-z,$
$\dot{z}$ $=15\;\text{sgn}(x_{3})|x_{3}|^{\frac{7}{10}}+66\;\text{sgn}(x_{2})|x_{2}|^{\frac{7}{13}}+80\;\text{sgn}(x_{1})|x_{1}|^{\frac{7}{16}}.$ (18)
The control input $u$ is designed as,
@@
Figure 2: LABEL:ado3_states, LABEL:ado3_ss and LABEL:ado3_d_tilde show the ultimate boundedness of state trajectories, sliding variable and disturbance estimation error, respectively, on the application of adaptive ADO-based FOITSMC. Moreover, the minimum estimated adaptive gain required for GUUB is shown in LABEL:ado3_k_hat.
@@
Adaptive super-twisting control (STC) strategies [17, 22, 25, 26] often use signum functions in adaptation laws, which can lead to over- or under-estimated gains. This occurs when gains increase despite a decreasing sliding variable magnitude, or vice versa. In contrast, the adaptive law in (11) is devoid of any such discontinuous functions, thus striving $\hat{k}$ to achieve the desired performance with lesser gain values.
Simulation results were generated using MATLAB R2021a’s ODE4 (Runge-Kutta) solver with a fixed step size of 0.001. Figures LABEL:ado3_states and LABEL:ado3_ss show that the states of (17) and the sliding variable, respectively, remain within specified bounds in finite time under the control input (19). Figure LABEL:ado3_u illustrates a chatter-free control input. The estimated $k$ is non-monotonic (as shown in fig. LABEL:ado3_k_hat) and it decreases (or increases) depending on whether the magnitude of the sliding variable decreases (or increases).
V Performance Comparison
@@
such that $|\dot{a}_{1}(x,t)|<\delta<\infty$, $\delta$ unknown. For any given arbitrary initial conditions $x(0)$, $\sigma(0)$, there exists a finite $t_{f}>0$ such that a real 22-sliding mode has been established $\forall\;t>t_{f}$ via control law
@@
$\dot{\alpha}$ $=\begin{cases}\omega_{1}\sqrt{\frac{\gamma_{1}}{2}}\text{sgn}\left(|\sigma|-\mu\right),\;\;&\alpha>\alpha_{m}\\ \eta,\;&\alpha\leqslant\alpha_{m}\end{cases}$ (22)
@@
Figure 3: The comparison result of the proposed algorithm with the existing ASTW in the presence of disturbances $\sin(2\pi t)$ and $\sin(2\pi t)+\texttt{ramp}(t)$, where $\texttt{ramp}(t)=0.5t\left(\text{sgn(t)}+1\right)$. LABEL:com_khat1 and LABEL:com_khat2 shows the comparison of estimated $k$ for both the cases. LABEL:com_ss1 and LABEL:com_ss2 illustrates the comparison in sliding variable for two kind of disturbances.
Simulations were conducted using MATLAB R2021a’s ODE4 (Runge-Kutta) solver with a fixed step size of 0.001. The disturbances $\sin(2\pi t)$ and $\sin(2\pi t)+\text{ramp}(t)$, where $\text{ramp}(t)=0.5t(\text{sgn}(t)+1)$, were deliberately chosen to test robustness against bounded and unbounded (with bounded time-derivative) disturbances. Figures LABEL:com_khat1 and LABEL:com_khat2 show that the proposed algorithm’s estimated gain $\hat{k}$ remains lower than that of ASTW in both cases. Figures LABEL:com_ss1 and LABEL:com_ss2 demonstrate that the sliding variable $s$ in the proposed algorithm stays closer to zero, in comparison to the existing ASTW, signifies greater robustness against the acting disturbances.
VI Case study
Attitude control and stabilisation of a rigid spacecraft under external disturbances have long been a focus of control research [30, 31, 32, 33, 34, 35] due to the system’s highly nonlinear dynamics, which pose significant challenges for control design. This section applies the proposed control algorithm to achieve robust attitude stabilisation of a rigid spacecraft. The unit-quaternion-based attitude dynamics of a rigid spacecraft are described by:
$\begin{cases}\dot{q}_{0}&=-\frac{1}{2}q_{v}^{T}\Omega,\\ \dot{q}_{v}&=\frac{1}{2}(q_{0}I_{3}+q_{v}^{\times})\Omega,\\ J\dot{\Omega}&=-\Omega^{\times}J\Omega+u+d,\end{cases}$ (24)
where the unit quaternion $q=[q_{0},q_{1},q_{2},q_{3}]^{T}=[q_{0},q_{v}]^{T}$ satisfies $q_{0}^{2}+q_{v}^{T}q_{v}=1$, with $q_{v}\in\mathbb{R}^{3}$ as the vector part and $q_{0}\in\mathbb{R}$ as the scalar component. Here, $\Omega\in\mathbb{R}^{3}$ is the angular velocity, $u\in\mathbb{R}^{3}$ is the control torque, $J\in\mathbb{R}^{3\times 3}$ is the symmetric positive definite inertia matrix, $I_{3}$ is the $3\times 3$ identity matrix, and the skew-symmetric operator for any vector $a=[a_{1},a_{2},a_{3}]^{T}$ is:
$a^{\times}=\begin{pmatrix}0&-a_{3}&a_{2}\\ a_{3}&0&-a_{1}\\ -a_{2}&a_{1}&0\end{pmatrix}$. (25)
@@ VI Case study
The exogenous matched disturbance $d$ is continuously differentiable such that each components $d_{i}$ of $d$ satisfy, $\dot{d}_{i}\leqslant k_{i},\;\;\text{for}\;i=1,2,3$ where $k_{i}$s are finite.
@@
[…]
@@
$k_{i}\hat{k}_{i}-\hat{k}_{i}^{2}$ $=-\left(\frac{k_{i}}{\sqrt{2}}-\frac{\hat{k}_{i}}{\sqrt{2}}\right)^{2}-\frac{\hat{k}_{i}^{2}}{2}+\frac{k_{i}^{2}}{2}\leqslant-\frac{\tilde{k}_{i}^{2}}{2}+\frac{k_{i}^{2}}{2}.$ (28)
@@
[…]
@@
Figure 4: This figure illustrates the performance of the proposed adaptive controller under matched disturbances. Sub-figures (LABEL:spacecraft_quaternions)–(LABEL:spacecraft_d_tilde) confirm the global boundedness of the attitude quaternions, angular velocities ($\Omega$), the sliding variable, and the disturbance estimation error. The corresponding continuous control input and the evolution of the adaptive gain ($\hat{k}$) are presented in (LABEL:spacecraft_u) and (LABEL:spacecraft_k_hat), respectively.
@@
The simulation results shown in fig. LABEL:spacecraft_quaternions and LABEL:spacecraft_omega, utilising the ODE 4 (Runge-Kutta) solver of MATLAB R2021a with an assigned fixed step size of 0.001, depict that the state trajectories remain close to the origin, bounded within a specified limit. A continuous control input has been employed using an adaptive ADO-based FOITSMC algorithm, as shown in Fig. LABEL:spacecraft_u, to ensure the global boundedness of error in the estimation of conditionally known disturbance and unknown gain $k$.